A baseball player hits a homerun over the left field fence, which is 104 m from home plate. The ball is hit at a point 1.00 m directly above home plate, with an intial velocity of 30 degrees above the horizontal. By what distance does the baseball clear the 3.00m high fence, if it passes over it 3.00 seconds after being hit.
Answer #1

let vo is the initial velocity of base ball.

in horizontal direction,

time taken to cross the fense, t = x/vox

3 = 104/(vo*cos(30))

vo = 104/(3*cos(30))

= 40.03 m/s

vertical height of the base ball when it crosses the fence,

y = yo + voy*t - (1/2)*g*t^2

= 1 + 40.03*sin(30)*3 - (1/2)*9.8*3^2

= 17.0 m

distance between the fense and base ball = 17.0 - 3.0

= 14.0 m <<<------Answer